# reference.bib

### v2

@@ 4 ConfusionPrompt: User-Side Models
Following the overall design and the requirements in privacy model, this section introduces how to enable ConfusionPrompt from the perspective of designing user-side models, including decomposer, generator, and recomposer.
@@ 4.1.1 Complexity Analysis
Theorem 4.1 (Complexity for Single Paragraph).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
@@
Theorem 4.2 (Complexity for Decomposed Prompt).
Let =$\mathbb{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompt with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
@@
Theorem 4.1 and 4.2 compares the complexities between single prompt and decomposed sub-prompt, where the proof is defered to appendix B. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. See Figure 4 for intuitive explanation.
@@ 4.1.2 Requirements of Ideal Decomposer
(1) A fundamental requirement for the decomposition process is to follow the collectively exhaustive principle [rasiel1999mckinsey]. In other words, the collection of responses to decomposed prompt should recover the whole response.
(2) An ideal decomposer would optimize the complexity, i.e., the required number of pseudo prompts, in two aspects: (i) each sub-prompt contain as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
@@ 4.1.3 Model Training
Following the spirit of Reinforcement Learning from Human/AI Feedback (RLHF/RLAIF) [christiano2017deep, lee2023rlaif], we train a decomposer that optimizes the aforementioned two objectives simultaneously. For training efficiency and stability, the reward model training and reinforcement learning stages are replaced with Direct Preference Optimization (DPO), which directly optimizes a model on the preference data [rafailov2023direct]. Therefore, the training prosedure of decomposer is as follows:
Supervised fine-tuning (SFT): We finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to decompose given a full prompt.
Preference data construction: We use the SFT model to generate various decompositions given a full prompt $p$ to generate a various decompositions $(p_{1},p_{2},...,p_{l})\sim\pi_{\mathrm{SFT}}(\cdot|p)$. The comparison data considering the two requirements/dimensions are collected by either human labelers or AI feedback, following the rules in section D.1.
DPO optimization: Given a full prompt $p$, we denote $p_{w}$ and $p_{d}$ as the preferred and dispreferred decomposition. Then the optimization objective is formulated as:
$\begin{gathered}\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{SFT}})=-\mathbb{E}_{(p,p_{w},p_{d})}\\ \left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(p_{w}|p)}{\pi_{\mathrm{SFT}}(p_{w}|p)}-\beta\log\frac{\pi_{\theta}(p_{d}|p)}{\pi_{\mathrm{SFT}}(p_{d}|p)}\right)\right]\end{gathered}$ (9)
where $\sigma$ is the logistic function, $\beta$ is a hyperparameter controlling the deviation from SFT model, and $\theta$ is the model parameters to optimize.
@@ 4.2 Generator
Requirements. The requirements for a pseudo prompt suggest the following criteria for an ideal generator: (1) The generator should produce pseudo prompt with fake information that’s semantically irrelevant to the correspondent attributes in genuine prompt. (2) The pseudo prompt should not be classified as fabricated prompt by a strong discriminator $D$.
Model training. The generator is trained using a similar procedure described in Section 4.1.3: (i) SFT: fine-tune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. (ii) Preference data construction: collect comparison data considering both semantic relevance and genuineness; see details in section D.2). (iii) DPO optimization: optimize the model on the comparison data.
@@ 4.3 Recomposer
The local re-composition LM combines the retrived sub-prompts and sub-responses to produce the final whole response. We train a local recomposer on a collection of demonstration data with supervise learning. Specifically, given a complete prompt $p$ and a set of sub-prompt-response pairs $\left((p_{1},r_{1}),...(p_{k},r_{k})\right)$, the recomposer $G_{r}$ is trained to maximize $\log p_{G_{r}}\left(r|p,(p_{1},r_{1}),...(p_{k},r_{k})\right)$ where $r$ is the final response.
@@ 4.4 Algorithm
Algorithm 1 outlines the procedure for our private inference framework ConfusionPrompt. The client starts with decomposing the full prompt into sub-prompts, and generate a collection of qualified fake prompts that meets the privacy requirement. It can be derived immediately from the algorithm that our protocol guarantees $(\lambda,\mu,\rho)$-privacy.
@@ 5.1 Experiment Setup
We evaluate our framework on two multi-hop reasoning datasets: StrategyQA [geva2021did] and MuSiQue [trivedi2022musique]. We consider ConfusionPrompt with different online LLMs including GPT-4, GPT-4-Turbo, and GPT-4o [openai2023gpt4]. We consider the privacy budget $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to section F.2 for the training of decomposer, generator, and recomposer.
@@ 5.2 Empirical Privacy Evaluation
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [vashisth2020gender] to predict the gender based on the user’s review.
@@ 5.3.1 Privacy Experiments
Figure 2 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$, genuineness $\rho$, similarity $\lambda$. It can be observed that: (1) decreasing significance $\mu$ consistently reduces the attack accuracy. (1) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more fluent. (3) There’s no clear relationship between the attack accuracy and similarity threshold $\lambda$. Such privacy parameter could be more related to the attribute inference attack discussed later.
@@
The results of attribute inference attack are detailed in F.4. A key observation is that when $1/\mu\geq 15$, the attack accuracies for ConfusionPrompt decrease to a plateau and is similar to that of Text2Text with $\epsilon\leq 1$.
@@ 5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [yang2018hotpotqa]. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against five baseline methods: (1) a local Llama2 [touvron2023llama] model with 7B parameters; (2) a local Vicuna [chiang2023vicuna] model with 13B parameters; (3) Text-to-text privatization (Text2Text) [feyisetan2020privacy], where the query is privatized with LDP by replacing each word with the perturbed token embeddings; (4) Paraphraser [utpala2023locally, mattern2022limits] that paraphrases the prompt with LDP mechanism using a language model; (5) querying online GPTs with no privacy protection. In the former two methods, the user downloads the open-source models and directly extract the final response from the model. In methods (3) and (4), the queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server. For the final approach, the complete prompt is transmitted to the service provider without any confusion strategy.
Table 1: Utility comparisons between our proposed ConfusionPrompt and baselines. ConfusionPrompt strikes the best balance between privacy preservation and utility.
@@
The results are reported in Table 1. From the results, we see that (1) online proprietary models outperform open-source models by over 11% and 24%, repectively, for StrategyQA and MuSiQue, indicating the strong motivation for users to use online service while preserving privacy. (2) ConfusionPrompt based on various online models consistently and significantly outperforms LDP-based methods, showing the effectiveness of our proposed method. Notably, our framework possesses a distinct advantage over perturbation-based methods: increasing the privacy budget does not adversely affect performance since the true queries are always provided to the server.
@@ 5.3.3 Decomposition Benefits
Theorem 4.1 and 4.2 show that the decomposition module could reduce the number of pseudo prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
@@
Figure 3 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth before decomposition, whereas it exhibits approximate linearity after decomposition. Our protocol offer substantial savings ranging from an average reduction of 1.9×1.9$\times$1.9 × at $\mu=1/10$ to an average reduction of 18.3×18.3$\times$18.3 × at $\mu=1/50$.

### v4

@@ 4 ConfusionPrompt: User-Side Models
Following the overall design and the requirements in privacy model, this section introduces how to enable ConfusionPrompt through the design of user-side models, including decomposer, generator, and recomposer.
@@ 4.1.1 Complexity Analysis
Theorem 11 (Complexity for Single Paragraph).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the genuine user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a single paragraph $|\boldsymbol{p}_{i}|=1$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
@@
Theorem 12 (Complexity for Decomposed Prompt).
Let $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ be a group of prompts, where $\boldsymbol{p}_{0}$ is the user prompt and the remaining ones are dummy prompts. Suppose each prompt in the group represents a sequence of decomposed prompt with the same intention $\boldsymbol{p}_{i}=[p_{i1},p_{i2},...,p_{il}]$, $\forall i\in[1,n]$. To achieve $(\lambda,\mu,\rho)$-privacy, it requires that
@@
Theorem 11 and 12 compare the complexities between single prompt and decomposed sub-prompt, where the proof is defered to Appendix A. In the ideal scenario, the complexity can be reduced to $\mathcal{O}\left(|U(\boldsymbol{p}_{0})|/\mu\right)$ after decomposition. An intuitive explanation is given in Figure 2.
@@ 4.1.2 Requirements of Ideal Decomposer
  • 1.
    A fundamental requirement for the decomposition process is to follow the collectively exhaustive principle [35]. In other words, the collection of responses to decomposed sub-prompts should recover the whole response.
  • 2.
    An ideal decomposer would optimize the complexity, i.e., the required number of pseudo prompts, in two aspects: (i) each sub-prompt contains as few attributes as possible; (ii) each attribute appears in as few prompts as possible.
@@ 4.1.3 Model Training
We design a two-stage training approach to optimize the above objectives. In the first stage, we finetune a pretrained language model (LM) on demonstration data for decomposition task, where the LM is trained to generate decomposition given a full prompt. In the second stage, the LM is prompted to output multiple decompositions for each input. From the generations, a subset of preferred decompositions is selected to further fine-tune the model. The procedure to select preferred examples are as follows:
  • 1.
    Evaluate whether each decomposition adheres to the MECE principle, discarding any that do not meet this standard.
  • 2.
    Among the MECE-compliant decompositions, assess the complexity of each using a predefined parameter $\mu$, and select the top $k$ examples for each prompt with the lowest complexity.
@@ 4.2 Generator
4.2.1 Requirements
The requirements for a pseudo prompt suggest the following criteria for an ideal generator:
  • 1.
    The generator should produce pseudo prompt with fake information that is semantically irrelevant to the correspondent attributes in genuine prompt.
  • 2.
    The pseudo prompt should not be classified as fabricated prompt by a strong discriminator $D$.
4.2.2 Model training.
The generator is trained using a similar procedure described in Section 4.1.3. In the first stage, we fine-tune a pretrained LM using demonstration data for replacement task, where the LM is trained to generate new prompt with replaced attributes. In the second stage, we collect the preferred generations considering both semantic relevance and genuineness, with criteria given as follows:
  • 1.
    Compute $s$, the similarity between the pseudo and original prompt.
  • 2.
    Compute $f$, the genuineness score for the pseudo prompt.
  • 3.
  • 4.
    Select the top $k$ examples for each prompt with the highest overall score.
@@ 4.3 Recomposer
The local re-composition LM combines the retrieved sub-prompts and sub-responses to produce the final whole response. We train a local recomposer on a collection of demonstration data with supervise learning. Specifically, given a complete prompt $p$ and a set of sub-prompt-response pairs $\left((p_{1},r_{1}),...(p_{k},r_{k})\right)$, the recomposer $G_{r}$ is trained to maximize $\log p_{G_{r}}\left(r|p,(p_{1},r_{1}),...(p_{k},r_{k})\right)$ where $r$ is the final response.
Figure 2: Example of decomposition savings in query complexity. Decomposition module reduces the query complexity from 9 to 6 under privacy requirement $\mu=3$. This reduction becomes more pronounced with larger values of $\mu$. For instance, the complexity decreases from 100 to 20 under privacy requirement $\mu=10$.
@@ 4.4 Algorithm
Algorithm 1 outlines the procedure for our private inference framework ConfusionPrompt. The client starts with decomposing the full prompt into sub-prompts and then generates a collection of qualified fake prompts that meet the privacy requirement. It can be derived immediately from the algorithm that our protocol meets $(\lambda,\mu,\rho)$-privacy.
Algorithm 1 ConfusionPrompt
2:Output: Final response $r$.
4:for $p_{0h}\in\boldsymbol{p}_{0}$ do
5:     User keeps sampling from the generator until there are $n$ distinct pseudo prompts with $n$ given by 9, where each one $1\leq i\leq n$ has prompt similarity $\mathrm{Sim}(p_{0h},p_{ih})\leq\lambda$ and genuineness score $\mathrm{Genu}(p_{ih})\geq\rho$.
6:end for
7:User sends the group of genuine and pseudo prompts $\boldsymbol{P}=\{\boldsymbol{p}_{0},\boldsymbol{p}_{1},...,\boldsymbol{p}_{n}\}$ to the server.
8:Server returns a collection of responses $(\boldsymbol{r}_{0},...,\boldsymbol{r}_{n})$.
9:User retrieves the response corresponding to the genuine prompt $\boldsymbol{r}_{0}$.
@@ 5.1 Experiment Setup
We evaluate our framework on two multi-hop reasoning datasets: StrategyQA [36] and MuSiQue [37]. We integrate ConfusionPrompt with different online LLMs including GPT-4, GPT-4-Turbo, and GPT-4o [7]. We consider the privacy parameters $\lambda\in[0.5,0.8]$, $1/\mu\in[5,50]$, and $\rho\in[1,4]$ unless specified. Refer to Appendix C.2 for the training of decomposer, generator, and recomposer.
We extract the private attributes of StrategyQA and MuSiQue through the steps below (see Appendix C.1 for details):
  • 1.
    Step 1: sample construction. We adopt a semi-automated method to extract private attributes from each sample. First, we extract the entities using a combination of two NER methods: Spacy [38] and Flair [39]. Then we manually correct and supplement the private attributes for each query.
  • 2.
  • 3.
@@ 5.2 Empirical Privacy Evaluation
Attribute Inference Attack: an attack that infers the sensitive features of records from either a group of prompts, or differentially privatized inputs. We rely on the twitter text dataset [40] to predict the gender based on the user’s review.
@@ 5.3.1 Privacy Experiments
Figure 3 visualizes the attack accuracy for prompt identification attack under various combinations of significance $\mu$, genuineness $\rho$, similarity $\lambda$. It can be observed that: (a) decreasing significance $\mu$ consistently reduces the attack accuracy. (b) As the genuineness threshold $\rho$ increases from 1 to 4, the attack accuracy approaches that of random guessing. The result reveals that it becomes harder for the attacker to distinguish the true query as the fake prompts become more realistic. (c) There’s no clear relationship between the attack accuracy and similarity threshold $\lambda$. Such privacy parameter could be more related to the attribute inference attack discussed later.
@@
Figure 4 presents the attack accuracies of attribute inference attack for both ConfusionPrompt and LDP-based methods, from which we can make the following observations: (1) Paraphrase generally leads to higher attack accuracies compared to Text2text. This can be attributed to the fact that the semantic meanings of certain key words remain unchanged during the paraphrasing process. (2) The attack accuracies decrease to a plateau when $1/\mu\geq 15$ or $\epsilon\leq 1$. This suggests that beyond these thresholds, further increasing the privacy protection metrics does not significantly reduce the attack accuracy. (3) The attack accuracy for ConfusionPrompt with $1/\mu\geq 15$ is similar to that of Text2Text with $\epsilon\leq 1$, indicating certain alignment for the privacy protection between LDP-based methods and our framework. (4) Decreasing the similarity threhold $\lambda$ helps to reduce the attribute inference attack accuracies.
@@ 5.3.2 Utility Evaluation
For StrategyQA, we report the accuracy scores (ACC) and area under the roc curve (AUC) to assess the the classification task. For MuSiQue, we report the F1 score, ROUGE-L, and exact matching (EM) to evaluate the question answering task [41]. We benchmark our framework with $\mu=15$, $\rho=4$, and $\lambda=0.5$ against five baseline methods: (1) a local Llama2 [42] model with 7B parameters; (2) a local Vicuna [43] model with 13B parameters; (3) Text-to-text privatization (Text2Text) [30], where the query is privatized with LDP by replacing each word with the perturbed token embeddings; (4) Paraphraser [31, 32] that paraphrases the prompt with LDP mechanism using a language model; (5) querying online GPTs with no privacy protection. In the former two methods, the user downloads the open-source models and directly extract the final response from the model. In methods (3) and (4), the queries are privatized with LDP privacy budget $\epsilon=10$ before transmission to the server. For the final approach, the complete prompt is transmitted to the service provider without any confusion strategy
Table 1: Utility comparisons between our proposed ConfusionPrompt and baselines.
@@
   The privacy parameters for ConfusionPrompt are fixed at $\mu=15$, $\rho=4$, and $\lambda=0.5$. The privacy budgets for LDP-based methods are set to $\epsilon=10$ and $\delta=10^{-4}$.
The results are reported in Table 1. From the results, we can see that (1) online proprietary models outperform open-source models by over 11% and 24%, repectively, for StrategyQA and MuSiQue, indicating the strong motivation for users to use online service while preserving privacy. (2) ConfusionPrompt based on various online models consistently and significantly outperforms LDP-based methods, showing the effectiveness of our proposed method. Notably, our framework possesses a distinct advantage over perturbation-based methods: increasing the privacy protection level does not adversely affect performance since the true queries are always provided to the server.
@@ 5.3.3 Decomposition Benefits
Theorem 11 and 12 show that the decomposition module could reduce the number of pseudo prompts required to be sent to the server. In this section, we empirically demonstrate practical benefits by analyzing the monetary cost of our ConfusionPrompt framework. In particular, we compute the monetary ratio, the ratio of the monetary cost for using ConfusionPrompt compared to the direct query method, both before and after the application of the decomposition module.
@@
Figure 5 presents the query cost benefits achieved through our decomposition module. It can be observed that the complexity gains become increasingly pronounced as the level of privacy protection increases. This can be attributed to the fact that the complexity experiences polynomial growth before decomposition, whereas it exhibits approximate linearity after decomposition. Our protocol offers substantial savings ranging from an average reduction of 1.9×1.9$\times$1.9 × at $\mu=1/10$ to an average reduction of 18.3×18.3$\times$18.3 × at $\mu=1/50$.